Let R be the region bounded below by the graph of y=(x^2)+5 and above by the graphs of y=6x and y=12. Find the volume of the solid obtained by revolving R about the x-axis.
Answer #1

el R be the region bounded below by the graph of y=x+5 and above by the graphs of Y = 6* f y = 12. find the volume of the sol